Suppose you are inside the water in a swimming pool near an edge. A friend is standing on the edge. Do you find your friend taller or shorter than his usual height?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

consider an object placed outside water surface , if observer is in water then he will see the image of object formed after refration through water ....

let initially object is placed at a distance x from surface

 for refraction from rarer to denser medium  (u is the refractive index of medium)

u/v - 1/x =u-1/R                                      (refractive index of air is 1)

 R=infinity

u/v=1/x

v=ux              (u>1)

so v>u

thus the height of friend who is outside is greater than original ...